### v5

@@ 3 Identification
Now that we have characterized response-groups, we seek to identify the probabilities of the corresponding response-groups in the treatment model. Let $P(t|z)\equiv\Pr(T_{i}=t|Z_{i}=z)$ denote the generalized propensity score. Under strict, one-to-one targeting, the response-groups are easily enumerated.
Proposition 3 (Counting response-groups under strict, one-to-one targeting).
Suppose that $p$ treatment values are targeted and $q$ are not. Under Assumptions 1 to 5, the number of response-groups is $N\equiv(p+2q)\times 2^{p-1}$.
As the probabilities of the response-groups must sum to one, we have $(N-1)$ unknowns. The data gives us the generalized propensity scores $P(t|z)$ for $(t,z)\in\mathcal{T}\times\mathcal{Z}$. The adding-up constraints $\sum_{t\in\mathcal{T}}P(t|z)=1$ for each $z\in\mathcal{Z}$ reduce the count of independent data points to $(\lvert\mathcal{T}\rvert-1)\times\lvert\mathcal{Z}\rvert=(p+q-1)(p+1)$.
$\mathcal{T}$ $p$ $q$ Unknowns Equations Required
LATE ( {0,1} 1 1 2 2 0
Example 1 {0,1,…,$\lvert\mathcal{T}\rvert-1$} 1 $\lvert\mathcal{T}\rvert-1$ 2$(\lvert\mathcal{T}\rvert-1)$ 2$(\lvert\mathcal{T}\rvert-1)$ 0
Example 2 {0,1,2} 2 1 7 6 1
Table 3: Identifying the sizes of the response groups under strict, one-to-one targeting
Table 3 shows some values of the number of equations and the number of unknowns $(N-1)$ for three examples. The first row has $\lvert\mathcal{T}\rvert=\lvert\mathcal{Z}\rvert=2$; it generates the standard LATE case, where the response group consists of never-takers ($C_{00}$), compliers ($C_{01}$), and always-takers ($C_{11}$). The second row is another case of exact identification. The third row shows that one restriction is required to identify the sizes of the response-groups for the $3\times 3$ model. More generally, the degree of underidentification increases exponentially with the number of targeted treatments $p$. The probabilities of the different groups are linked to the generalized propensity scores by a system of linear equations. Under strict, one-to-one targeting, this system takes a simple form, as mentioned in Section 3.
Proposition 4 (Identifying equations for group sizes under strict, one-to-one targeting).
Under Assumptions 1 to 5, the generalized propensity scores satisfy the following system of equations, for all $(z,t)\in\mathcal{Z}\times\mathcal{T}$:
(3.1) $\displaystyle P(t|z)$ $\displaystyle=\sum_{A\subset\mathcal{T}^{\ast}\;\setminus\;\{z\}}{\rm 1\kern-3.99994pt1}(t\in\bar{A})\Pr(i\in C(A,t))$
$\displaystyle+\sum_{A\subset\mathcal{T}^{\ast}}{\rm 1\kern-3.99994pt1}(t=z\in A)\;\sum_{\tau\in\bar{A}}\Pr(i\in C(A,\tau))$
where we denote $\bar{A}\equiv(\mathcal{T}\,\setminus\,\mathcal{T}^{\ast})\operatorname*{{\bigcup}}A$.
While this may look cryptic, it is directly related to Corollary 1: the first line corresponds to $z\in\mathcal{T}\,\setminus\,A_{i}$ and $t=\underline{t}_{i}$, and the second line corresponds to $t=z\in A_{i}$. The set $\bar{A}_{i}\equiv(\mathcal{T}\,\setminus\,\mathcal{T}^{\ast})\operatorname*{{\bigcup}}A_{i}$ contains the non-targeted treatments and those for which $T_{i}(t)=t$.
To simplify the exposition, we introduce one more element of notation. For any $z\in\mathcal{Z}$ and $t\in\mathcal{T}$, we define the conditional average outcome by $\bar{E}_{z}(t)\equiv\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)|Z_{i}=z)$. For any response-group $C$ and treatment value $t\in\mathcal{T}$, we define the group average outcome as $\mathbb{E}(Y_{i}(t)|i\in C)$. While the conditional average outcomes $\bar{E}_{z}(t)$ are directly identified from the data, the group average outcomes of course are not. We do know that some of them are zero; and that they combine with the group probabilities to form the conditional average outcomes. We will repeatedly use the following identity from Heckman and Pinto (2018, Theorem T-1):
@@
Let $z\in\mathcal{Z}$ and $t\in\mathcal{T}$. Then
$\bar{E}_{z}(t)=\sum_{C=C_{R}\;|\;R(z)=t}\mathbb{E}(Y_{i}(t)\;|i\in C)\Pr(i\in C).$
In addition,
$\mathbb{E}(Y_{i}\;|Z_{i}=z)=\sum_{t\in\mathcal{T}}\bar{E}_{z}(t).$
Under strict, one-to-one targeting, the set of response-groups $C=C_{R}$ such that $R(z)=t$ is as enumerated in Proposition 4: it consists of
@@
  • all $C(A,t)$ such that $A\subset\mathcal{T}^{\ast}\;\setminus\;\{z\}$ and $t\in\bar{A}$;
  • and, if $t=z$, all $C(A,\tau)$ for $z\in A\subset\mathcal{T}^{\ast}$ and $\tau\in\bar{A}$.
The combination of Lemma 1 and of either Proposition 2 (under strict targeting) or Proposition 4 (under strict, one-to-one targeting) does not exhaust the empirical content of the model. A succession of papers [^17] See Balke and Pearl (1997), Kitagawa (2015), Mourifié and Wan (2017), Kédagni and Mourifié (2020) and Sun (2023). has given necessary and sometimes sufficient conditions for data to be rationalized under an instrument exclusion restriction. Most recently, Bai and Tabord-Meehan (2024) characterized the sharp testable implications of ARUM under joint independence [^18] That is, when $Z_{i}$ is independent of $\{(Y_{i}(t),T_{i}(z)):(t,z)\in\mathcal{T}\times\mathcal{Z}\}$..
To simplify the exposition, we will state our results in terms of effects of the treatment on the expectation of the outcomes; they hold, however, for any measurable function of the outcome $f(Y)$. Note that if we chose $f(Y)={\rm 1\kern-3.99994pt1}(Y\leq t)$ for some value $t$, we would identify the effects of the treatment on the cumulative distribution function of the outcome. By inversion, we would recover the quantile treatment effects.
3.1 Positive Selection
We will sometimes make use of an identifying assumption that we call positive selection. It obtains when for some treatment value $t$ and response groups $C\neq C^{\prime}$ that sometimes choose $t$, [^19] That is, for which $\Pr(T_{i}=t|i\in C)$ and $\Pr(T_{i}=t|i\in C^{\prime})$ are nonzero. we have $\mathbb{E}(Y_{i}(t)|i\in C)\leq\mathbb{E}(Y_{i}(t)|i\in C^{\prime})$. The identifying power of positive selection depends on the context; we will illustrate it in Corollaries 2 and 3 as well as in our application to Head Start in Section 4. A slightly different definition would replace $Y_{i}(t)$ with a treatment effect $Y_{i}(t)-Y_{i}(t^{\prime})$; we explore this variant in Corollary 4.
@@ 3.2 The Binary Instrument Model
Recall that with a binary instrument, strict targeting is trivially satisfied. Under one-to-one targeting, Proposition 4 can be applied directly to some of the rows of Table 3.
3.2.1 Identification Under One-to-one Targeting
The second row of Table 3 shows that the group probabilities are just identified in our Example 1 under strict, one-to-one targeting. Proposition 4 gives $2(T-1)$ independent equations: for $t\neq 1$,
@@ 3.2.1 Identification Under One-to-one Targeting
Proposition 5 (Response-group probabilities in Example 1 under one-to-one targeting).
Under Assumptions 1 to 5, the following probabilities are identified:
@@
Proposition 6 (Group average outcomes in Example 1 under one-to-one targeting).
Under Assumptions 1 to 5, the following group average outcomes are point-identified:
@@
Proposition 6 shows that we only identify a known convex combination of the $(\lvert\mathcal{T}\rvert-1)$ LATEs [^20] We use the term “LATEs” for the average treatment effects on the various complier groups. Throughout the remainder of the paper, we assume, as is standard, that probability differences appearing in the denominator of estimands are always nonzero.. This formula is reminiscent of Angrist and Imbens (1995, Theorem 1), which deals with a different model in which treatments are ordered. It is possible to re-derive our identification results in Propositions 5 and 6 using the general framework of Heckman and Pinto (2018). We provide details in Appendix D.
@@
3.2.2 Adding Identification Constraints
@@
First note that the probabilities of the response-groups can be identified under weaker restrictions than one-to-one targeting. Suppose for instance that $z=1$ targets all treatment values $t\geq 1$: we have $U_{1}(t)>U_{0}(t)$ for all $t\geq 1$. Then the complier groups $C_{t0}$ for $t\geq 1$ must be empty. To see this, suppose that $T_{i}(0)=t\geq 1$. This implies $U_{0}(t)+u_{it}>U_{0}(0)+u_{i0}=u_{i0}$. Adding up these inequalities gives $U_{1}(t)+u_{it}>u_{i0}$, and $T_{i}(1)$ cannot be 00 0.
@@
In this $2\times 3$ model under one-to-one targeting, there are five response-groups, as illustrated in Figure 3. Proposition 6 shows that the Wald estimator only identifies
@@
where $\alpha_{0}=(P(0|0)-P(0|1))/(P(1|1)-P(1|0))$ is point-identified. Corollary 2 shows that adding inequality (3.4) yields bounds on the corresponding LATEs.
[…]
@@
Corollary 2 (Positive selection and treatment effects in the $2\times 3$ model under one-to-one targeting).
@@
(3.5) $\displaystyle\mathbb{E}\left[Y_{i}(1)|i\in C_{01}\right]\leq\mathbb{E}\left[Y_{i}(1)|i\in C_{21}\right],$
then the local average treatment effects for $C_{01}$ and $C_{21}$ are partially identified:
(3.6) $\displaystyle\begin{split}\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)|i\in C_{01}\right]&\leq\frac{\bar{E}_{1}(1)-\bar{E}_{0}(1)}{P(1|1)-P(1|0)}-\frac{\bar{E}_{0}(0)-\bar{E}_{1}(0)}{P(0|0)-P(0|1)},\\ \mathbb{E}\left[Y_{i}(1)-Y_{i}(2)|i\in C_{21}\right]&\geq\frac{\bar{E}_{1}(1)-\bar{E}_{0}(1)}{P(1|1)-P(1|0)}-\frac{\bar{E}_{0}(2)-\bar{E}_{1}(2)}{P(2|0)-P(2|1)},\end{split}$
Moreover, (3.5) implies the following testable prediction:
(3.7) $\frac{\bar{E}_{0}(2)-\bar{E}_{1}(2)}{P(2|0)-P(2|1)}\geq\frac{\bar{E}_{0}(0)-\bar{E}_{1}(0)}{P(0|0)-P(0|1)}.$
If (3.5) holds at equality, then the two statements in (3.6) and the testable prediction in (3.7) also become equalities, and the two LATEs are point-identified.
The lower bounds on the local average treatment effects for $C_{01}$ and $C_{21}$ may not be sharp; on the other hand, they are easy to estimate from sample averages. It is a topic for future research to obtain the sharp bounds and develop a corresponding method for estimation and inference.

### v6

@@ 3 Identification
We use the standard counterfactual notation: $T_{i}(z)$ and $Y_{i}(t,z)$ denote respectively potential treatments and outcomes. Let ${\rm 1\kern-3.99994pt1}(A)$ denote the indicator of set $A$. The validity of the instruments requires the usual exclusion restriction together with appropriate independence conditions.
Assumption 5 (Valid Instruments).
  • (i)
    $Y_{i}(t,z)=Y_{i}(t)$ for all $(t,z)\in\mathcal{T}\times\mathcal{Z}$.
  • (ii)
    For all $(t,z)\in\mathcal{T}\times\mathcal{Z}$ and all response groups $C$,
    $\displaystyle\Pr(i\in C\mid Z_{i}=z)$ $\displaystyle=\Pr(i\in C);$
    $\displaystyle\mathbb{E}(Y_{i}(t)\mid i\in C,Z_{i}=z)$ $\displaystyle=\mathbb{E}(Y_{i}(t)\mid i\in C).$
One could impose a stronger condition than Assumption˜5(ii), such as the joint independence of $\{(Y_{i}(t),T_{i}(z)):(t,z)\in\mathcal{T}\times\mathcal{Z}\}$ from $Z_{i}$. However, we adopt the weaker mean-independence assumption for $Y_{i}(t)$ in Assumption˜5(ii), as it is sufficient for the results presented in this paper. Assumption˜5(ii) can be viewed as a generalization of Assumption 2 in Huber, Laffers, and Mellace (2017), who only consider the case of binary instruments and binary treatments. Under Assumption˜5, we define $T_{i}:=T_{i}(Z_{i})$ and $Y_{i}:=Y_{i}(T_{i})$. Throughout the paper, we assume that we observe $(Y_{i},T_{i},Z_{i})$ for each $i$.
To simplify the exposition, we introduce one more element of notation. For any $z\in\mathcal{Z}$ and $t\in\mathcal{T}$, we define the conditional average outcome by
$\bar{E}_{z}(t)\equiv\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)\mid Z_{i}=z).$
For any response-group $C$ and treatment value $t\in\mathcal{T}$, we define the group average outcome as $\mathbb{E}(Y_{i}(t)\mid i\in C)$. Our goal is to identify the proportions of each response-group, $\Pr(i\in C)$, and their group average outcomes.
While the conditional average outcome $\bar{E}_{z}(t)$ is directly identified from the data, the group average outcomes are not; they combine with the group probabilities to form the conditional average outcomes. We will repeatedly use the following identity:
@@
Let Assumption 5 hold. Then, for all $(t,z)\in\mathcal{T}\times\mathcal{Z}$,
$\displaystyle P(t|z)$ $\displaystyle=\sum_{C_{R}\,|\,R(z)=t}\Pr(i\in C_{R});$
$\displaystyle\bar{E}_{z}(t)$ $\displaystyle=\sum_{C_{R}\;|\;R(z)=t}\mathbb{E}(Y_{i}(t)\;|i\in C_{R})\Pr(i\in C_{R}).$
If there are $\lvert\mathcal{R}\rvert$ response-groups, the first equation in Lemma 1 is a system of $(\lvert\mathcal{T}\rvert-1)\times\lvert\mathcal{Z}\rvert$ equations with $(\lvert\mathcal{R}\rvert-1)$ unknowns. Under monotonicity, the binary-binary model $\lvert\mathcal{T}\rvert=\lvert\mathcal{Z}\rvert=2$ generates the standard LATE case, where the response-groups consist of never-takers ($C_{00}$), compliers ($C_{01}$), and always-takers ($C_{11}$). As is well known, the sizes of these three groups are just identified. On the other hand, even under Assumption˜4, one restriction is required to identify the sizes of the response-groups for the $3\times 3$ model. [^16] We prove this in Proposition 8. More generally, the degree of underidentification of group sizes tends to increase exponentially with the number of targeted treatments, like the number of response-groups.
Each of our assumptions on targeting reduces the number of response-groups and therefore the degree of underidentification. Take our strongest assumption: universal targeting. Then the set of response-groups $C=C_{R}$ such that $R(z)=t$ is as enumerated in Proposition 4: it consists of
@@
  • all $C(A,t)$ such that $A\subset\mathcal{T}^{\ast}\setminus\{z\}$ and $t\in\bar{A}\equiv A\operatorname*{{\bigcup}}(\mathcal{T}\setminus\mathcal{T}^{\ast})$;
  • if $t=z$, all $C(A,\tau)$ for $z\in A\subset\mathcal{T}^{\ast}$ and $\tau\in\bar{A}$.
This gives directly the system of identifying equations.
Proposition 4 (Identifying equations under universal targeting).
Let Assumptions 1, 2, 4, and 5 hold. Then, for all $(z,t)\in\mathcal{Z}\times\mathcal{T}$:
(3.1) $\begin{split}\bar{E}_{z}(t)&=\sum_{A\subset\mathcal{T}^{\ast}\setminus\{z\}}{\rm 1\kern-3.99994pt1}(t\in\bar{A})\mathbb{E}[Y_{i}(t)\mid i\in C(A,t)]\Pr(i\in C(A,t))\\ &\quad+\sum_{A\subset\mathcal{T}^{\ast}}{\rm 1\kern-3.99994pt1}(t=z\in A)\sum_{\tau\in\bar{A}}\mathbb{E}[Y_{i}(t)\mid i\in C(A,\tau)]\Pr(i\in C(A,\tau)),\\[10.00002pt] P(t|z)&=\sum_{A\subset\mathcal{T}^{\ast}\setminus\{z\}}{\rm 1\kern-3.99994pt1}(t\in\bar{A})\Pr(i\in C(A,t))\\ &\quad+\sum_{A\subset\mathcal{T}^{\ast}}{\rm 1\kern-3.99994pt1}(t=z\in A)\sum_{\tau\in\bar{A}}\Pr(i\in C(A,\tau)).\end{split}$
3.1 Positive Selection and the Empirical Content of the Model
We now introduce an identifying assumption that we call positive selection. It obtains when a function $h$ of the vector of potential outcomes $\bm{Y}_{i}\equiv(Y_{i}(0),\ldots,Y_{i}(\lvert\mathcal{T}\rvert-1))$ has a lower expectation for a response-group $C$ than for another response-group $C^{\prime}$:
(3.2) $\mathbb{E}(h(\bm{Y}_{i})\;|\;i\in C)\leq\mathbb{E}(h(\bm{Y}_{i})\;|\;i\in C^{\prime}).$
Two leading examples are (a) $h(\bm{Y}_{i})\equiv Y_{i}(t)$ for some $t$ and (b) $h(\bm{Y}_{i})\equiv Y_{i}(t)-Y_{i}(t^{\prime})$ for some $t\neq t^{\prime}$. In the $2\times 2$ model, Huber, Laffers, and Mellace (2017) consider a variance of (a), which they call mean dominance. The identifying power of positive selection depends on the context. We will illustrate it in the form of (a) in Proposition˜7, as well as in our application to Head Start in Section 4. We explore variant (b) in Proposition˜12.
Before turning to specific applications, let us discuss the characterization of the empirical content of our framework under positive selection. A series of papers [^17] See Balke and Pearl (1997), Kitagawa (2015), Mourifié and Wan (2017), Kédagni and Mourifié (2020), and Sun (2023). has provided necessary and, in some cases, sufficient conditions for data to be rationalized under an instrument exclusion restriction. Most recently, Bai, Huang, and Tabord-Meehan (2025) characterized the sharp testable implications of what they call “encouragement design” under joint independence, that is, when $Z_{i}$ is independent of $\{(Y_{i}(t),T_{i}(z)):(t,z)\in\mathcal{T}\times\mathcal{Z}\}$. Because we assume only mean independence in Assumption˜5 and impose positive selection via (3.2), their testable implications involving outcomes $Y_{i}$ do not directly apply to our framework. However, their sharp restrictions on the generalized propensity scores $P(t|z)$ do apply, and we discuss these in our leading examples below. It is worth noting that while they do not impose an ARUM structure, their “encouragement design” condition is satisfied in our context only under our strongest assumption: universal targeting.
In the following sections, we characterize the empirical content for the restrictions given in (3.1), together with a positive selection assumption $\mathbb{E}(Y_{i}(t)\mid i\in C)\leq\mathbb{E}(Y_{i}(t)\mid i\in C^{\prime})$ for given $C$ and $C^{\prime}$. The primitives of the model are the response-group probabilities $\Pr(i\in C)$ and the group average outcomes $\mathbb{E}(Y_{i}(t)\mid i\in C)$. The restrictions implied by (3.1) generate bilinear equality constraints for $\mathbb{E}(Y_{i}(t)\mid i\in C)$ and $\Pr(i\in C)$, while positive selection yields linear inequality constraints for $\mathbb{E}(Y_{i}(t)\mid i\in C)$. This makes it difficult to obtain simple yet general characterizations. Instead, we present explicit results for specific models in the next two subsections.
In particular, we study the identification region of local average treatment effects,
$\mathbb{E}[Y_{i}(t^{\prime})-Y_{i}(t)\mid i\in C],$
for specific pairs $(t,t^{\prime})$ and subgroups $C$.
@@ 3.2 The Binary Instrument Model
Recall that with a binary instrument, strict targeting is trivially satisfied.
3.2.1 Identification Under Universal Targeting
Under one-to-one targeting, $z=1$ targets only one instrument value, which we call $t=1$; and targeting is universal. Proposition˜4 can be applied directly and the group probabilities are just identified in our Example˜1. Proposition 4 gives $2(\lvert T\rvert-1)$ independent equations: for $t\neq 1$,
@@ 3.2.1 Identification Under Universal Targeting
Proposition 5 (Response-group probabilities in Example˜1 under universal targeting).
Under Assumptions 1, 2, 4, and 5, the following probabilities are identified:
@@
Proposition 6 (Group average outcomes in Example˜1 under universal targeting).
Under Assumptions 1, 2, 4, and 5, the following group average outcomes are point-identified:
@@
Proposition˜6 shows that we only identify a known convex combination of the $(\lvert\mathcal{T}\rvert-1)$ LATEs [^18] We use the term “LATEs” for the average treatment effects on the various complier groups. Throughout the remainder of the paper, we assume, as is standard, that probability differences appearing in the denominator of estimands are always nonzero.. This formula is reminiscent of Angrist and Imbens (1995, Theorem 1), which deals with a different model in which treatments are ordered. It is possible to re-derive our identification results in Propositions 5 and 6 using the general framework of Heckman and Pinto (2018). We provide details in Appendix˜E.
@@
3.2.2 Alternative Identification Constraints
@@
First note that the probabilities of the response-groups can be identified under weaker restrictions than one-to-one targeting. Suppose for instance that $z=1$ targets all treatment values $t\geq 1$: we have $U_{1}(t)>U_{0}(t)$ for all $t\geq 1$. Then the complier groups $C_{t0}$ for $t\geq 1$ must be empty. To see this, suppose that $T_{i}(0)=t\geq 1$. This implies $U_{0}(t)+u_{it}>U_{0}(0)+u_{i0}=u_{i0}$. Adding up these inequalities gives $U_{1}(t)+u_{it}>u_{i0}$, and $T_{i}(1)$ cannot be $0$.
@@
Truncated Moment Bounds with Bounded Outcomes
One alternative approach is to use truncated moment bounds, a method initiated by Horowitz and Manski (1995) and popularized by Lee (2009) in the context of randomized experiments with sample attrition. See also Huber, Laffers, and Mellace (2017) for extensions to the IV setting with binary instruments and binary treatments, and Semenova (2025) for refinements of Lee bounds. However, to obtain informative bounds within this framework, it is often necessary to assume that the support of the outcome variable is bounded.
Monotone Treatment Responses.
It is sometimes natural to assume that treatments are ordered and that potential outcomes are weakly increasing in the treatment level, that is, $Y_{i}(t)\geq Y_{i}(t^{\prime})$ whenever $t\geq t^{\prime}$, as in Manski (1997). A weaker variant imposes monotonicity only in expectation within a response group $C$, $\mathbb{E}[Y_{i}(t)\mid i\in C]\geq\mathbb{E}[Y_{i}(t^{\prime})\mid i\in C]$ for all $t\geq t^{\prime}$, as in Marx (2024) [^19] See Section 4.2.2 of the job market paper version of Marx (2024), available at https://www.dropbox.com/scl/fi/dq5hnkwwbsjpluuonpcwf/marx-jmp.pdf?rlkey=1zpwmnx5dgtmre5urp5hdzof5&e=4&dl=0. . These restrictions differ conceptually from our positive-selection assumptions. Monotone treatment response (MTR) constrains outcomes across treatment levels within a fixed group, whereas positive selection compares average outcomes across response groups at a given treatment level. Since our analysis focuses on unordered treatments, we focus here on selection-based restrictions.
@@
In this $2\times 3$ model under universal targeting, there are five response-groups, as illustrated in Figure˜3. Proposition 6 shows that the Wald estimator only identifies
@@
where $\alpha_{0}=(P(0|0)-P(0|1))/(P(1|1)-P(1|0))$ is point-identified. Proposition˜7 shows that adding inequality (3.5) yields bounds on the corresponding LATEs.
$u_{i1}-u_{i0}$$u_{i2}-u_{i0}$$C_{01}$$C_{00}$$C_{22}$$C_{11}$$C_{21}$$P_{0}$$P_{1}$
@@
Proposition 7 (Positive selection and treatment effects in the $2\times 3$ model under universal targeting).
@@
(3.6) $\displaystyle\mathbb{E}\left[Y_{i}(1)\mid i\in C_{01}\right]\leq\mathbb{E}\left[Y_{i}(1)\mid i\in C_{21}\right],$
then the sharp identified region for the local average treatment effects for $C_{01}$ and $C_{21}$
$L_{01}:=\mathbb{E}\!\left[Y_{i}(1)-Y_{i}(0)\mid i\in C_{01}\right]\;\text{ and }\;L_{21}:=\mathbb{E}\!\left[Y_{i}(1)-Y_{i}(2)\mid i\in C_{21}\right]$
is the half-line in the $(L_{01},L_{21})$ plane with equation
(3.7) $L_{01}\{P(0\mid 0)-P(0\mid 1)\}+L_{21}\{P(2\mid 0)-P(2\mid 1)\}\\ =(\bar{E}_{1}(1)-\bar{E}_{0}(1))-(\bar{E}_{0}(0)-\bar{E}_{1}(0))-(\bar{E}_{0}(2)-\bar{E}_{1}(2))$
limited by
(3.8) $\displaystyle\begin{split}L_{01}&\leq\frac{\bar{E}_{1}(1)-\bar{E}_{0}(1)}{P(1\mid 1)-P(1\mid 0)}-\frac{\bar{E}_{0}(0)-\bar{E}_{1}(0)}{P(0\mid 0)-P(0\mid 1)},\\ L_{21}&\geq\frac{\bar{E}_{1}(1)-\bar{E}_{0}(1)}{P(1\mid 1)-P(1\mid 0)}-\frac{\bar{E}_{0}(2)-\bar{E}_{1}(2)}{P(2\mid 0)-P(2\mid 1)}.\end{split}$
If (3.6) holds with equality, then the two inequalities in (3.8) hold at equality as well, and $L_{01}$ and $L_{21}$ are point identified.
The lower bounds on the local average treatment effects for $C_{01}$ and $C_{21}$ can be estimated directly from sample averages and are sharp in the sense that they incorporate all model restrictions. Specifically, the characterization relies only on two types of constraints: (i) the linear equality constraint in (3.7), which corresponds to the special case of the second equality in (3.4) from Proposition˜6, and (ii) the linear inequality constraint implied by positive selection in (3.6). Accordingly, the lower bounds in (3.8) constitute the explicit solutions to the associated linear programming problems. This explicit sharp characterization is possible because the response-group probabilities are point identified, as established in Proposition˜5.